OpenStudy (anonymous):

Let {q1,q2,...,qn} be the orthonormal vectors obtain after applying the Gram-Schmidt procedure on an mxn matrix A. Show that A = QR, where Q is an mxn with columns q1,q2,...,qn and R is an nxn upper triangular matrix.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

QR decomposition eh, i wouldnt be able to do this on the spot. I would have to read up on it to fill in the parts ive forgotten over theyears :)

OpenStudy (anonymous):

yes..QR decomposition.. its okay..thank you :)

OpenStudy (amistre64):

Q = [e1, e2, e3, ..., en] R = [a1.e1; a2.e1; a3.e1; ...; an.e1 ] [ 0 ; a2.e2; a3.e2; ...; an.e2 ] [ 0 ; 0 ; a3.e3; ...; an.e2 ] [ ........................................... ] [ 0 ; 0 ; 0 ; ...; an.en ]

OpenStudy (amistre64):

Show that: ak = [e1; e2; e3;...; ek; ....;en] * [ak.e1; ak.e2; ak.e3; ...; ak.ek; 0; 0; ....; 0]

OpenStudy (amistre64):

well, since everything after ak.ek is 0; we can effectively ignore it. So demonstrate that: ak = [e1; e2; e3;...; ek] * [ak.e1; ak.e2; ak.e3; ...; ak.ek]

OpenStudy (amistre64):

the dot product of ak.er is equal to the rth element of ak giving us a vector of say: [ak1; ak2; ak3; ak4; ...; akk] does that make sense?

OpenStudy (amistre64):

my Q row is off, row1 is equal to e1: 1,0,0,0,...,0 row2 is equal to e2: 0,1,0,0,...,0 etc

OpenStudy (amistre64):

ak = [ek] * [ak.e1; ak.e2; ak.e3; ...; ak.ek] ak = [ek] * [ak1; ak2; ak3; ...; akk] ak = the kth element of ak but this might be way too much work to show and they might be expecting you to use thrms that i have long forgotten

OpenStudy (anonymous):

okay..thanks a lot :)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@radar

OpenStudy (radar):

Sorry, this is above my pay-grade. Haven't a clue.

OpenStudy (anonymous):

its okay..thank you @radar :)

OpenStudy (phi):

This is a pain to explain. call the columns of A= a1, a2, etc The idea is to use the definition of the Gram-Schmidt procedure to show that q1 is a1 (normalized) , a2 = a1 - a2 projected onto a1, a3 = a3 - a3 proj a1 - a3 proj a2 Set up the equation A = Q R the columns of R will select out the columns of Q to combine a1 (first col of A) = q1 scaled by |a1|, and no other columns. i.e. the first column of R will have 1 entry at 1,1, and zeros below it. a2 will be a linear combination of q1 and q2 only... if you follow this pattern you find R is upper triangular.

OpenStudy (radar):

I knew this was above my paygrade.